OpenStudy (anonymous):

How to factor? 12x^6-27y^4. I know to take out the three, but I don't know what to do from there. Any help? Thanks.

OpenStudy (anonymous):

once you take out the 3, you have the difference of two squares

OpenStudy (anonymous):

\[3(4x^6-9y^4)\] the inner piece is like \(a^2-b^2=(a+b)(a-b)\) with \(a=2x^3, b=3y^2\)

OpenStudy (anonymous):

Ah, okay, I didn't know the formula. Thank you.

OpenStudy (anonymous):

May I ask for the formula if it is a^2+b^2, @satellite73 ?

jimthompson5910 (jim_thompson5910):

`a^2+b^2` is a sum of squares and can't be factored (assuming 'a' and 'b' have nothing in common)

OpenStudy (anonymous):

I'm working on the problem now, just gimme a minute and we'll see if I got it right. :-)

OpenStudy (anonymous):

I got 2x^5+3y^2)(2x^4-3y^2)

OpenStudy (anonymous):

what @jim_thompson5910 said you can factor the "difference of two squares" but not the "sum of two squares"

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

@satellite73 How did you get 2x^3?

OpenStudy (anonymous):

the exponent for the \(x\) term should be \(3\) in both cases

OpenStudy (anonymous):

that is a good question

OpenStudy (anonymous):

i visualized \(4x^6\) as \((2x^3)^2\)

OpenStudy (anonymous):

I thought you just subtract. Since to the squared is to, you'd minus two from 6.

OpenStudy (anonymous):

So you divide the exponent?

jimthompson5910 (jim_thompson5910):

satellite73 is using the rule \[\LARGE (x^a)^b = x^{a*b}\] notice how the inner exponent 'a' is being multiplied with the outer exponent b

jimthompson5910 (jim_thompson5910):

so with 4x^6, satellite73 broke up the 6 into 3*2 and used the rule above

OpenStudy (anonymous):

Does my understanding of it? You divide?

OpenStudy (anonymous):

*does it work ,too?

jimthompson5910 (jim_thompson5910):

yeah 6/2 = 3

jimthompson5910 (jim_thompson5910):

and 4/2 = 2

OpenStudy (anonymous):

Does that work in general or was I just lucky there?

jimthompson5910 (jim_thompson5910):

well again, notice how 6 factors into 3*2 we want the *2 portion so we can ultimately have something squared so that's why we divide by 2 to figure out that other factor

jimthompson5910 (jim_thompson5910):

if we had say x^10, we'd break it up like 5*2 how do we find 5? well we can say 10/2 = 5

OpenStudy (anonymous):

Ah, okay, I am working on another problem right now. It would mean so much if you could stay for a bit to see if I got it right. :-)

OpenStudy (anonymous):

This may take more for more as me as it involves big numbers. Thanks for your patience.

OpenStudy (anonymous):

Finally got the factors. 28 and 28.

OpenStudy (anonymous):

Hold on a bit.

OpenStudy (anonymous):

(4x^2+7)^3

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

I just realized I didn't give you the problem. XD

OpenStudy (anonymous):

16x^4+56x^2+49

OpenStudy (anonymous):

your answer is correct

OpenStudy (anonymous):

Actually, according to the calculator, it's squared not cubed. :(

OpenStudy (anonymous):

oh no it isn't the exponent should be 2

OpenStudy (anonymous):

\[16x^4+56x^2+49\] is a "perfect square"

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\] in your case \(a=4x^2,b=7\)

OpenStudy (anonymous):

Thank you. I'm fine for now, but I may be good. Thanks again for your time! :-)